# Segmenting Action-Value Functions Over Time-Scales in SARSA via TD($\Delta$)

### v2

@@ 6.2 OpenAI Gym Environments
The following plots show the results of two algorithms, k-step SARSA($\Delta$) and k-step SARSA TD, on the FrozenLake-v1, Taxi-v3, CliffWalking-v0, and Blackjack environments from OpenAI Gym. The y-axis shows the average rewards obtained by the algorithms over episodes, and the x-axis shows the number of episodes needed to train the algorithms. Episodes range from 0 to 100,000 episodes. The shaded area surrounding each curve shows the standard error over 10 random seeds. The algorithms both use k= 32, which is the number of steps. The remaining hyperparameter settings are displayed in the table 1.
Hyperparameter Value
seed (10) np.random.randint(40, 45)
Learning Rate ($\alpha$) [0.1, 0.3, 0.5, 0.7, 1.0]
Discount Factor ($\eta$) [0.992, 0.95, 0.996, 0.99, 1.0]
k-Step [1, 4, 8, 16, 32, 128, 256]
Number of Episodes 100000
Max Epsilon 0.4
The red line denotes K-step SARSA ($\Delta$), while the blue line represents K-step SARSA TD in both Fig. 4(a) and Fig. 4(b). Fig. 4(a) illustrates the FrozenLake-v1 environment, whereas Fig. 4(b) depicts the Taxi-v3 environment. Fig. 4(a) demonstrates that k-step SARSA ($\Delta$) continually attains superior average rewards relative to k-SARSA TD as training advances. K-step SARSA ($\Delta$) demonstrates considerable enhancement in average rewards over the initial 20,000 episodes, signifying accelerated learning, but k-step SARSA TD exhibits lesser increase in average rewards in comparison to k-step SARSA ($\Delta$) during the same episodes. Following 40,000 episodes, the efficacy of k-step SARSA($\Delta$) stabilizes while consistently surpassing that of k-step SARSA TD, indicating enduring reliability. k-step SARSA regularly exhibits inferior performance relative to k-step SARSA ($\Delta$).
Fig. 4(b) illustrates that K-Step SARSA ($\Delta$) regularly achieves superior rewards (i.e., nearer to -200) during the training phase, signifying enhanced learning efficiency and overall performance. Fig. 4(b) also demonstrates that K-Step SARSA ($\Delta$) stabilizes in the initial phases, indicating robustness and the capacity to converge to an effective policy, whereas K-Step SARSA TD commences at considerably lower rewards (approximately -1800) and exhibits negligible improvement, remaining predominantly flat throughout the training process. The K-Step SARSA TD algorithm encounters difficulties in acquiring an effective policy in this setting compared to K-Step SARSA ($\Delta$). Experiment result show that K-Step SARSA ($\Delta$) significantly outperforms K-Step SARSA TD in the Taxi-v3 environment, achieving much higher average rewards and faster convergence.
(a) FrozenLake-v1
(b) Taxi-v3
@@
The Fig. 5(a) and 5(b) illustrates the learning effectiveness of two variations of the K-Step SARSA algorithm, namely K-Step SARSA ($\Delta$) and K-Step SARSA TD, across 100,000 episodes in the CliffWalking-v0 and Blackjack-v1 environments. Both algorithms are assessed using k=32, a hyperparameter that specifies the number of steps in the updating process. The trend of average rewards in fig. 5(a) demonstrates that both algorithms show an increasing pattern in average rewards as the number of episodes rises, indicating improved policy performance with continued learning. Nevertheless, K-Step SARSA ($\Delta$) often outperforms K-Step SARSA TD, achieving higher average rewards over the episodes.
The red line in fig. 5(a) denoting K-Step SARSA ($\Delta$) demonstrates a considerable enhancement in the initial episodes (about 20,000 episodes), indicating a swifter convergence to optimal policies relative to K-Step SARSA TD. In the 100,000th episode, K-Step SARSA ($\Delta$) attains a reward near -1050, whereas K-Step SARSA TD obtains a slightly lower payout of approximately -1120. This demonstrates the enhanced efficacy of K-Step SARSA ($\Delta$) in this task.
@@
The average reward of approximately +0.15 is dependably achieved across all episodes by the red curve in fig. 5(b) that represents K-Step SARSA ($\Delta$). The blue curve in fig. 5(b), which represents K-Step SARSA TD, suggests stability; however, it presents an average reward that is significantly lower, at approximately -0.2, across all episodes. K-Step SARSA ($\Delta$) markedly surpasses K-Step SARSA TD regarding average rewards. The favorable average reward of K-Step SARSA ($\Delta$) demonstrates its superior suitability for policy optimization in the Blackjack-v1 environment, whereas the negative reward for K-Step SARSA TD signifies inferior performance.
@@
Finally, when comparing the results of all four environments, the findings demonstrate the constant advantage of K-Step SARSA ($\Delta$) over K-Step SARSA TD across all evaluated environments. This superiority is seen through elevated average rewards, accelerated convergence, and enhanced stability in learning outcomes. Environments characterized by deterministic state transitions, such as CliffWalking-v0 and FrozenLake-v1, demonstrate the clear superiority of K-Step SARSA ($\Delta$), but intricate stochastic environments like Blackjack-v1 and Taxi-v3 further highlight its resilience. The performance graph of deterministic environments, such as CliffWalking-v0 and FrozenLake-v1, resembles an exponential growth plateau curve due to predictable state transitions. In contrast, the performance graph of complex stochastic environments, like Blackjack-v1 and Taxi-v3, appears as a flat line because of the uncertainty and variability in state transitions.
7 Conclusion
With the help of TD($\Delta$), SARSA($\Delta$) improves SARSA by integrating multiple discount factors, leading to better stability and performance. With this approach, we can get a more precise breakdown of action-values while still enjoying the benefits of on-policy learning. In addition to the deterministic and stochastic OpenAI Gym environments, we also conducted trials in a tabular synthetic environment, specifically the five-star ring MDP. In all investigated environments, our experimental results demonstrate that SARSA($\Delta$), our suggested approach, outperforms k-step SARSA TD in terms of greater average rewards, faster convergence, and more stable learning trajectories. The results of this work advance our understanding of multi-step RL algorithms and how they are applied to various sequential decision-making tasks. This information can aid in the creation of robust algorithms for unpredictable, real-world environments in the future. In essence, the Q-Learning version of TD($\Delta$) extends this technique. We decided to take care of it later.
@@ Acknowledgments
The proof utilizes induction. Initially, it is necessary to establish that the base case is valid at n=0. This statement is evidently valid given our initial assumption, particularly in the context of zero initialization. At a specific time-step n, we assume the validity of the statement, namely, $\sum_{m=0}^{M}\theta_{n}^{m}=\theta_{n}^{\eta}$. We will now demonstrate that it also holds at the subsequent time step, n+1.
$\sum_{m=0}^{M}\theta_{n+1}^{m}$ $=\sum_{m=0}^{M}\bigg{(}\theta_{n}^{m}+\alpha_{m}\bigg{(}G_{n}^{M,\lambda_{m}}-\hat{D}_{m}(s_{n},a_{n})\bigg{)}\phi(s_{n},a_{n})\bigg{)}$
@@
$=\theta_{n}^{\eta}+\sum_{m=0}^{M}\alpha_{m}\bigg{(}\sum_{k=n}^{\infty}(\lambda_{m}\eta_{m})^{k-n}\delta_{k}^{m}\bigg{)}\phi(s_{n},a_{n})\text{ , Assumption based on induction.}$.
$=\theta_{n}^{\eta}+\alpha\sum_{k=n}^{\infty}(\lambda\eta)^{k-n}\underbrace{\bigg{(}\sum_{m=0}^{M}\delta_{k}^{m}\bigg{)}}_{*}\phi(s_{n},a_{n})\text{ , Because of $\alpha_{m}=\alpha$, $\lambda_{m}\eta_{m}=\lambda_{\eta}$,$\forall m$}$
To show that $\sum_{m=0}^{M}\theta_{n+1}^{m}=\theta_{n+1}^{\eta}$, we must establish that the term (*) = $\sum_{m=0}^{M}\delta_{k}^{m}$ is equivalent to the conventional TD error $\delta_{k}^{\eta}$.
$\sum_{m=0}^{M}\delta_{k}^{m}$ $=r_{k}+\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})-\hat{Q}_{\eta_{0}}(s_{k},a_{k})+\sum_{m=1}^{M}\bigg{(}(\eta_{m}-\eta_{m-1})\sum_{q=0}^{m-1}\langle\theta_{n}^{q},\phi(s_{k+1},a_{k+1})\rangle+\eta_{m}\langle\theta_{n}^{m},\phi(s_{k+1},a_{k+1})\rangle-\langle\theta_{n}^{m},\phi(s_{k},a_{k})\rangle\bigg{)}$
$=r_{k}+\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})+\bigg{\langle}\sum_{m=1}^{M}\bigg{(}\eta_{m}\sum_{q=0}^{m}\theta_{n}^{q}-\eta_{m-1}\sum_{q=0}^{m-1}\theta_{n}^{q}\bigg{)},\phi(s_{k+1},a_{k+1})\bigg{\rangle}-\bigg{\langle}\sum_{m=0}^{M}\theta_{n}^{m},\phi(s_{k},a_{k})\bigg{\rangle}$
$=r_{k}+\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})+\eta_{m}\bigg{\langle}\sum_{m=0}^{M}\theta_{n}^{m},\phi(s_{k+1},a_{k+1})\bigg{\rangle}-\eta_{0}\bigg{\langle}\theta_{n}^{0},\phi(s_{k+1},a_{k+1})\bigg{\rangle}-Q_{\eta}(s_{k},a_{k})$
$=r_{k}+\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})+\eta\hat{Q}_{\eta}(s_{k+1},a_{k+1})-\eta_{0}\hat{Q}_{\eta_{0}}(s_{k+1},a_{k+1})-\hat{Q}_{\eta}(s_{k},a_{k})$
$=r_{k}+\eta\hat{Q}_{\eta}(s_{k+1},a_{k+1})-\hat{Q}_{\eta}(s_{k},a_{k})$
$=\delta_{k}^{\eta}$ (32)
@@
The separated action-value function is denoted by D in this form, and the update is given in terms of the operator $\mathrm{T}_{\lambda}$, which is responsible for applying time-scale separation. In the event that $10\leq\lambda\eta<$ 1, the formula is considered to be well-defined. This is because the spectral norm of the operator $\lambda\eta P$ is less than 1, and thus, the Eq. $I-\lambda\eta P$ is invertible. On the other hand, when the value of $\lambda$ is bigger than one, the equivalence between Eqs. 34 and 35 is lost. This is not a problem because, in fact, we use the TD error for training, which, according to our expectations, corresponds to the definition of $\mathrm{T}_{\lambda}$ that is provided in Eq. 35. Now, let’s have a look at the contraction property that the operator $\mathrm{T}_{\lambda}$ possesses. To begin, the Eq. designated as 35 can be rewritten as follows:
@@
Here the expectations are over a random trajectory under $\pi$; thus $\mathbf{E}[r_{\iota}](\iota\leq k\mathit{-}1)$ denotes the expected value of the $\iota$th reward received, while $\mathbf{E}[Q^{\pi}(s_{k},a_{k})]$ is the expected value of the true value function at the $k$th state-action reached. The phased TD(k) update sums the terms $\eta^{\iota}\bigg{(}\frac{1}{n}\bigg{)}\sum_{i=1}^{n}r_{\iota}^{i}$, whose expectations are exactly the $\eta^{\iota}\mathbf{E}[r_{\iota}]$ appearing above [9].
[…]
@@
If we assume n and the probability of exceeding an $\epsilon$ value is fixed to be atmost $\delta$, we can solve for the resulting value of $\epsilon$:
@@
So by Hoeffding, if we have n samples, with probability at least $1\mathit{-}\delta$,
@@
Since we are dealing with k different state-action pairs, we use a union bound. To ensure the probability holds for all k hypotheses, adjust $\delta$ to $\delta/k$:
@@
We can now assume that all $\mathbb{E}[r_{i}]$ terms are estimated to at least $\epsilon$ accuracy. Substituting this back into the definition of the k-step TD update we get
@@
and then the second term is bounded by $\Delta^{\hat{Q}_{\eta}}_{n-1}$ by assumption. Hence, combining these terms, we get the overall bound for SARSA:
@@
$\hat{D}_{m,n}(s,a)=\frac{1}{n}\sum_{x-1}^{n}\bigg{(}\sum_{i=1}^{k_{m}-1}(\eta_{m}^{i}-\eta_{m-1}^{i})r_{i}^{(x)}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})\hat{Q}_{\eta_{m-1}}(s_{k}^{(x)},a_{k}^{(x)})+\eta_{m}^{k_{m}}\hat{D}_{m}(s_{k}^{(x)},a_{k}^{(x)})\bigg{)}$ (44)
We know that according to the multi-step update rule 17 for $m\geq 1$:
@@
$\hat{D}_{m,n}(s,a)-D_{m}(s_{n},a_{n})=\sum_{i=1}^{k_{m}-1}\bigg{(}\eta_{m}^{i}-\eta_{m-1}^{i}\bigg{)}\bigg{(}\frac{1}{n}\sum_{x-1}^{n}r_{i}^{(x)}-\mathbb{E}[r_{i}]\bigg{)}+\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\bigg{(}\sum_{q=0}^{m-1}\hat{D}_{q}(s_{k}^{(x)},a_{k}^{(x)})-\mathbb{E}[D_{m}(s_{k},a_{k})]\bigg{)}$
$+\eta_{m}^{k_{m}}\bigg{(}D_{m}(s_{k}^{(x)},a_{k}^{(x)})-\mathbb{E}[D_{m}(s_{k},a_{k})]\bigg{)}$ (46)
Assume that $k_{0}\leq k_{1}\leq...k_{m}=k$, the D estimates share at most $k_{m}=k$ reward terms $\frac{1}{n}\sum_{x-1}^{n}r_{i}^{(x)}$, Using Hoeffding inequality and union bound, we obtain that with probability $1-\delta$, each k empirical average reward $\frac{1}{n}\sum_{x-1}^{n}r_{i}^{(x)}$ deviates from the true expected reward $\mathbb{E}[r_{i}]$ by at most $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$. Hence, with probability $1-\delta$, $\forall m\geq 1$, we have:

### v4

@@ 6.2 OpenAI Gym Environments
(a) FrozenLake-v1.
(b) Taxi-v3.
(c) Blackjack-v1.
(d) CliffWalking-v0.
Figure 3: OpenAI Gym control used for all experiments. In order from left to right: Frozen Lake, Taxi, Blackjack, and Cliff Walking.
Two algorithms, k-step SARSA($\Delta$) and k-step SARSA TD, were tested in the FrozenLake-v1, Taxi-v3, CliffWalking-v0, and Blackjack environments of OpenAI Gym. The results are shown in the figures below. The y-axis displays how many awards the algorithms got on average across episodes, while the x-axis shows how many episodes it took to train the algorithms. Episodes range from 0 to 100,000 episodes. The shaded area around each curve displays the standard error for 10 random seeds. Both algorithms utilize k = 32, which is the number of steps.
The red line denotes K-step SARSA ($\Delta$), while the blue line represents K-step SARSA TD in both Fig. 4a and Fig. 4b. Fig. 4a presents the FrozenLake-v1 environment, while Fig. 4b showcases the Taxi-v3 environment. Fig. 4a illustrates that k-step SARSA ($\Delta$) consistently achieves higher average rewards compared to k-SARSA TD as training progresses. Over the first 20,000 episodes, K-step SARSA ($\Delta$) shows fast learning with a significant rise in average rewards. However, k-step SARSA TD displays a smaller increase in average rewards when compared to k-step SARSA ($\Delta$) during the same episodes. After 40,000 episodes, k-step SARSA ($\Delta$) performance levels off and regularly beats k-step SARSA TD, showing that it is reliable over the long-term. k-step SARSA consistently performs worse than k-step SARSA ($\Delta$).
Fig. 4b illustrates that K-Step SARSA ($\Delta$) regularly achieves superior rewards (i.e., nearer to -200) during the training phase, signifying enhanced learning efficiency and overall performance. Fig. 4b also demonstrates that K-Step SARSA ($\Delta$) stabilizes in the initial phases, indicating robustness and the capacity to converge to an effective policy, whereas K-Step SARSA TD commences at considerably lower rewards (approximately -1800) and exhibits negligible improvement, remaining predominantly flat throughout the training process. The K-Step SARSA TD algorithm encounters difficulties in formulating an effective policy in this context relative to K-Step SARSA ($\Delta$). The results of the experiment indicate that K-Step SARSA ($\Delta$) demonstrates a significant advantage over K-Step SARSA TD in the Taxi-v3 environment, attaining considerably higher average rewards and achieving faster convergence.
(a) FrozenLake-v1.
(b) Taxi-v3.
@@
Figs. 5a and 5b demonstrate the learning effectiveness of two variations of the K-Step SARSA algorithm: K-Step SARSA ($\Delta$) and K-Step SARSA TD. This analysis is based on 100,000 episodes conducted in the CliffWalking-v0 and Blackjack-v1 environments. The evaluation of both algorithms is conducted with k=32, which is a hyperparameter indicating the number of steps involved in the updating process. The overall pattern of average rewards shown in Fig. 5a indicates that both algorithms exhibit an upward trajectory in average rewards as the number of episodes increases, reflecting enhanced policy performance through ongoing learning. Nonetheless, K-Step SARSA ($\Delta$) routinely exhibits better performance than K-Step SARSA TD, leading to increased average rewards across the episodes.
In comparison to K-Step SARSA TD, the red line in Fig. 5a representing K-Step SARSA ($\Delta$) shows a significant improvement in the first episodes (about 20,000 episodes), suggesting a faster convergence to optimum policies. K-Step SARSA ($\Delta$) receives a reward close to -1050 in the 100,000th episode, whereas K-Step SARSA TD receives a somewhat smaller reward of around -1120. This illustrates how K-Step SARSA ($\Delta$) performs better on this task.
@@
The average reward of approximately +0.15 is dependably achieved across all episodes by the red curve in Fig. 5b that represents K-Step SARSA ($\Delta$). The blue curve in Fig. 5b, which represents K-Step SARSA TD, suggests stability; however, it presents an average reward that is significantly lower, at approximately -0.2, across all episodes. K-Step SARSA ($\Delta$) markedly surpasses K-Step SARSA TD regarding average rewards. The favorable average reward of K-Step SARSA ($\Delta$) demonstrates its superior suitability for policy optimization in the Blackjack-v1 environment, whereas the negative reward for K-Step SARSA TD signifies inferior performance.
@@
In conclusion, the comparison of results across all four environments reveals a consistent advantage of K-Step SARSA ($\Delta$) over K-Step SARSA TD in each case. This advantage is demonstrated by increased average rewards, quicker convergence, and enhanced stability in learning results. Environments defined by deterministic state transitions, such as CliffWalking-v0 and FrozenLake-v1, clearly showcase the advantages of K-Step SARSA ($\Delta$). In contrast, more complex stochastic environments like Blackjack-v1 and Taxi-v3 further emphasize its robustness. The performance graph of deterministic environments, such as CliffWalking-v0 and FrozenLake-v1, resembles an exponential growth plateau curve due to predictable state transitions. On the contrary, the performance graph for sophisticated stochastic environments, such as Blackjack-v1 and Taxi-v3, resembles a flat line due to the uncertainty and unpredictability in state transitions.
Table 2 presents the average rewards along with confidence intervals (CI) for each algorithm within their respective environments. Kruskal statistical tests [45] were conducted at a significance level of $\alpha=0.05$ to compare K-step SARSA TD and K-Step SARSA ($\Delta$) baseline models.
Table 2: This table presents a comparative analysis of K-Step SARSA TD and K-Step SARSA ($\Delta$) across multiple OpenAI Gym Toy Text environments.
Algorithm Environments
Blackjack-v1 CliffWalking-v0 FrozenLake-v1 Taxi-v3
[…]
6.3 OpenAI Gym Atari Environments
To assess the efficacy of the proposed PPO-TD($\lambda$, SARSA($\Delta$)) method, experiments were conducted in two Atari environments: SeaquestNoFrameskip-v4 and QbertNoFrameskip-v4 [3], and the results were compared with those obtained from the standard PPO algorithm [39]. The results shown in Figs. 7a and  7b show that in both environments, the proposed technique achieves better stability and better long-term rewards. The experiments tracked average rewards during training episodes ranging from 10,000 to 100,000. The experiments employ the code cited in Kostrikov[44].
(a) SeaquestNoFrameskip-v4
(b) QbertNoFrameskip-v4
Figure 6: The OpenAI Gym Atari control framework is utilized for conducting long-horizon experiments. Arranged from left to right: (a) SeaquestNoFrameskip-v4 and (b) QbertNoFrameskip-v4.
In SeaquestNoFrameskip-v4, PPO-TD($\lambda$, SARSA($\Delta$)) demonstrates superior performance compared to PPO-Standard at four out of five evaluation points, as illustrated in Fig. 7a. Following 100K episodes, the mean reward for PPO-TD($\lambda$, SARSA($\Delta$)) is around 14.35, while PPO-Standard attains 14.25, reflecting a 0.7% enhancement. PPO-Standard has a pronounced peak after 60K episodes, achieving an average reward of around 15.25, after which it declines and plateaus. In contrast, PPO-TD($\lambda$, SARSA($\Delta$)) exhibits more refined learning dynamics and sustains stable performance following its peak at 40K episodes.
The performance difference in QbertNoFrameskip-v4 is more substantial, as illustrated in Fig. 7b. At 100K episodes, PPO-TD($\lambda$, SARSA($\Delta$)) attains an average reward of approximately 275, whereas PPO-Standard achieves about 210, leading to an improvement of approximately 31%. It is noteworthy that PPO-TD($\lambda$, SARSA($\Delta$)) starts with lower rewards, around 155 at 10K episodes, in contrast to 210 for PPO-Standard. However, it exhibits a steady progress, finally surpassing the baseline after 60K episodes, and reaching a high of around 290 episodes after 80K episodes. The wider shaded region for the proposed technique in the earlier phases indicates more variance, which diminishes as training progresses—signifying improved convergence and learning efficiency with time.
In conclusion, these results highlight the advantages of integrating PPO and SARSA($\Delta$) with TD($\lambda$, $\Delta$). The suggested method enhances the stability of action-value estimates across various time scales and fosters superior long-term performance in dense-reward environments, especially in intricate games such as Qbert.
(a) SeaquestNoFrameskip-v4
(b) QbertNoFrameskip-v4
Figure 7: Comparison of performance between PPO-Standard and PPO-TD($\lambda$, SARSA($\Delta$)) across two Atari games: (a) SeaquestNoFrameskip-v4 and (b) QbertNoFrameskip-v4. The x-axis denotes the quantity of training episodes (measured in thousands), whereas the y-axis illustrates the average rewards obtained. The shaded areas represent the standard deviation calculated from 10 runs with different random seeds.
Table 3 shows the average rewards for each algorithm in their own environments, along with confidence intervals (CI). We used Kruskal statistical tests [45] at a significant level ($\alpha=0.05$) to compare the PPO-TD($\lambda$, SARSA($\Delta$))/PPO baseline models.
Table 3: Comparative analysis of PPO and PPO-TD($\lambda$, SARSA($\Delta$)) across various OpenAI Gym Atari environments.
Algorithm Atari Environments
Seaquest Qbert
[…]
7 Discussion and Conclusions
By combining SARSA with TD($\Delta$) and various discount factors, SARSA($\Delta$) improves SARSA’s stability and performance. This method maintains the benefits of on-policy learning while enabling a more precise breakdown of action-values. To evaluate the effectiveness of SARSA($\Delta$), experiments were performed in both deterministic and stochastic environments from OpenAI Gym, as well as in a synthetic tabular environment—specifically, the five-star ring MDP. Furthermore, enhancements in performance within dense-reward environments of the Atari suite were demonstrated through a PPO-based variant of SARSA($\Delta$). The experimental results across all examined environments indicate that the proposed method, SARSA($\Delta$), surpasses k-step SARSA TD by achieving higher average rewards, quicker convergence, and more consistent learning trajectories. The PPO-based variant of SARSA ($\Delta$) also demonstrates superior performance compared to standard PPO algorithms in dense-reward Atari environments, as evidenced by the experimental results outlined in Section 6.3. This research contributes to our knowledge of multi-step RL algorithms and its applicability to diverse sequential decision-making tasks.
Limitations: The proposed SARSA($\Delta$) method has four limitations that require attention: (i) Memory Overhead: When using a large number of $\Delta$ components, the memory needs increase since the component-level updates scale linearly with the number of discount factors. This technique requires an extra O(M) in storage compared to conventional SARSA. (ii) Stochastic Environments: Best results may not be attained through rigid scheduling of $\eta_{m}$ in scenarios characterized by high unpredictability of rewards. In certain cases, adaptive scheduling of $\eta_{m}$ can boost both the stability and the effectiveness of learning. (iii) Partial Observability Markov Decision Process (POMDP): Current formulations are inapplicable to environments characterized by partial observability, as they presuppose full observability of the Markov Decision Process (MDP). (iv) Extension to off-policy: There are numerous issues associated with adapting SARSA($\Delta$) to off-policy algorithms such as Q-learning. Off-policy learning has the potential to significantly increase variance in long-horizon components, which may subsequently lead to instability during the training process.
Future Directions: (i) Incorporation with Off-Policy Q-Learning: The adaptability of the SARSA($\Delta$) idea might be enhanced by applying it to off-policy algorithms, similar to the Q-learning version of TD($\Delta$). (ii) Applications to Multi-Agent Systems: In a multi-agent system, the agents coordinate on aligned subgoals by sharing $D_{m}$ components. (iii) Adaptive Scheduling via Meta-Gradients: Performance in various environments may be improved by using meta-gradient approaches for dynamically adjusting $\eta_{m}$. (iv) Extension to Partially Observable: Adapting SARSA($\Delta$) for environments characterized by partial observability is an effective approach to enhance its relevance in real-world applications. We decided to take care of it later.
@@ Acknowledgments
This research is funded by the Innovation Teams of Ordinary Universities in Guangdong Province under Grants (2021KCXTD038, 2023KCXTD022), the Key Laboratory of Ordinary Universities in Guangdong Province (2022KSYS003), Key Discipline Research Ability Improvement Project of Guangdong Province (2021ZDJS043, 2022ZDJS068), the Natural Science Foundation of Guangdong Province (2022A1515010990), the Research Fund of the Department of Education of Guangdong Province Grants (2022KTSCX079, 2023ZDZX1013, 2022ZDZX3012, 2022ZDZX3011, 2023ZDZX2038), Chaozhou Engineering Technology Research Center (z25025) and Hanshan Normal University project (XY202105). We express our gratitude to the editors and referees for their invaluable suggestions and remarks.
@@
M.H. participated in all stages, from Conceptualization design to writing—original draft preparation. G.Z., and X.D. funding acquisition. X.C. and Y.Z. supervision and project administration. W.H., L.M., S.Q., Z.Z., and P.W. investigation. Z.U. and N.J. validation and formal analysis.
@@
The corresponding authors may be contacted to request data related to this paper.
@@ Appendix A PROOFS
The separated action-value function is denoted by D in this form, and the update is given in terms of the operator $\mathrm{T}_{\lambda}$, which is responsible for applying time-scale separation. In the event that $0\leq\lambda\eta<1$, the formula is considered to be well-defined. This is because the spectral norm of the operator $\lambda\eta P$ is less than 1, and thus, Eq. $I-\lambda\eta P$ is invertible. On the other hand, when the value of $\lambda$ is bigger than one, the equivalence between Eqs. 31 and 32 is lost. This is not a problem because, in fact, training is based on the TD error, which, according to expectations, corresponds to the definition of $\mathrm{T}_{\lambda}$ that is provided in Eq. 32. Now, let’s have a look at the contraction property that the operator $\mathrm{T}_{\lambda}$ possesses. To begin, the Eq. designated as 32 can be rewritten as follows:
@@
In SARSA, the $\mathit{Q-value}$ function, Q(s, a), is estimated using samples of state-action pairs and their corresponding rewards. Let $Q_{n}(s,a)$ denote the estimate of the $\mathit{Q-value}$ at time n. For $\mathit{n}$ samples, by Hoeffding’s inequality provides a guarantee for a variable that is bounded between [1,+1][-1,+1], that, the probability that the sample mean deviates from the true mean by more than $\epsilon$ is expressed as follows:
@@
$\mathbb{P}\bigg{(}\bigg{|}\frac{1}{n}\sum_{x=1}^{n}r_{i}^{(x)}\mathit{-}\mathbb{E}[r_{i}]\bigg{|}\geq\epsilon\bigg{)}\leq 2e^{\bigg{(}\mathit{-}\frac{2n^{{}^{2}}\epsilon^{2}}{\sum_{x=1}^{n}(b\mathit{-}a)^{2}}\bigg{)}}\textit{where a=-1 and b=+1}$
@@
Assuming that n and the probability of surpassing an $\epsilon$ value is fixed to be atmost $\delta$, we can derive the corresponding value of $\epsilon$:
@@
So by Hoeffding’s inequality, for n samples, the following holds with probability at least $1\mathit{-}\delta$,
@@
Since the analysis involves k different state-action pairs, a union bound is applied. To ensure the probability holds for all k hypotheses, adjust $\delta$ to $\delta/k$:
@@
It is now assumed that all $\mathbb{E}[r_{i}]$ terms are estimated to at least $\epsilon$ accuracy. Substituting this back into the definition of the k-step TD update we get
@@
and then the second term is bounded by $\Delta^{\hat{Q}_{\eta}}_{n-1}$ by assumption. Therefore, the combination of these terms results in the comprehensive bound for SARSA:
@@
$\begin{split}\hat{D}_{m,n}(s,a)=\frac{1}{n}\sum_{x-1}^{n}\biggl{(}\sum_{i=1}^{k_{m}-1}(\eta_{m}^{i}-\eta_{m-1}^{i})r_{i}^{(x)}+(\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}})\hat{Q}_{\eta_{m-1}}(s_{k}^{(x)},a_{k}^{(x)})+\eta_{m}^{k_{m}}\hat{D}_{m}(s_{k}^{(x)},a_{k}^{(x)})\biggr{)}\end{split}$ (41)
According to the multi-step update rule 16 for $m\geq 1$:
@@
$\begin{split}\hat{D}_{m,n}(s,a)-D_{m}(s_{n},a_{n})&=\sum_{i=1}^{k_{m}-1}\bigg{(}\eta_{m}^{i}-\eta_{m-1}^{i}\bigg{)}\bigg{(}\frac{1}{n}\sum_{x-1}^{n}r_{i}^{(x)}-\mathbb{E}[r_{i}]\bigg{)}+\bigg{(}\eta_{m}^{k_{m}}-\eta_{m-1}^{k_{m}}\bigg{)}\\ &\bigg{(}\sum_{q=0}^{m-1}\hat{D}_{q}(s_{k}^{(x)},a_{k}^{(x)})-\mathbb{E}[D_{m}(s_{k},a_{k})]\bigg{)}+\eta_{m}^{k_{m}}\bigg{(}D_{m}(s_{k}^{(x)},a_{k}^{(x)})-\mathbb{E}[D_{m}(s_{k},a_{k})]\bigg{)}\end{split}$
Assume that $k_{0}\leq k_{1}\leq...k_{m}=k$, the D estimates share at most $k_{m}=k$ reward terms $\frac{1}{n}\sum_{x-1}^{n}r_{i}^{(x)}$, Using Hoeffding inequality and union bound, it follows that, with probability $1-\delta$, each k empirical average reward $\frac{1}{n}\sum_{x-1}^{n}r_{i}^{(x)}$ deviates from the true expected reward $\mathbb{E}[r_{i}]$ by at most $\epsilon=\sqrt{\frac{2log(2k/\delta)}{n}}$. Hence, with probability $1-\delta$, $\forall m\geq 1$, the following holds:
@@
Appendix B Hyperparameters
The experiments primarily utilize the hyperparameters recommended by Romoff 771 et al.[1], Schulman et al.[39] and Kostrikov[44]. The experiments utilize the following randomly generated seeds: 125125, 513, 90135, 81212, 3523401, 15709, 17, 0, 8412, and 1153780. Additional hyperparameter configurations are detailed in Table 4.
Hyperparameter Value
seed (10) np.random.randint(40, 45)
Learning Rate ($\alpha$) [0.1, 0.3, 0.5, 0.7, 1.0]
Discount Factor ($\eta$) [0.992, 0.95, 0.996, 0.99, 1.0]
k-Step [1, 4, 8, 16, 32, 128, 256]
Number of Episodes 100000
Max Epsilon 0.4
Clipping parameter 0.1
Number of actors 8
Horizon(T) 129
Number of Epochs 3
Lambda ($\lambda$) 0.95